enzh-CNfrdeeliwitjakorues
CTD Processing Summary CalCOFI 1701RL CTD Preliminary Data
Download 1701RL CTD raw cast files zipped Download 1701RL Preliminary CTD + bottle data 

General CTD Notes - data acquisition notes, logistics, processing - see below.
Please note that these regressions are generated from preliminary CTD vs bottle data and will be reprocessed once final bottle data are available. CTD temperatures and salinities do not usually change but oxygen, estimated chlorophyll-a, estimated nitrate may change significantly after point-checking. Questionable or mistrip bottle data are removed from these comparisons but may be visible on the CTD.csv plots. For this cruise and future cruises, both primary & secondary sensor profiles vs bottle data will be generated and archived in the downloadable CTD+Bottle data files. These plots are under the "csv-plots\Primary" & "csv-plots\Secondary" subdirectories.

CTD sensor corrections derived by comparing CTD sensor data, 4sec average prior to bottle closure, to bottle samples
Dual T, S, & O2 Primary Sensor Secondary Sensor
Salinity offset (bottle - CTD salinity; > 350m only; Seabird SBE4) -0.0004 -0.0034
Oxygen ml/L (dual Seabird SBE43; 2° O2 Calib Off) y = 1.0091x-0.0197
R² = 0.9995
y = 1.0559x + 0.0019
R² = 0.9995
Oxygen umol/Kg (dual Seabird SBE43; 2° O2 Calib Off) y =0.0106x+0.9593
R² = 0.9995
y = 1.0575x + 0.1822
R² = 0.9995
Single sensors Linear Polynomial
Nitrate - ISUS 4sec ave voltage vs Bottle NO3 (Frank's Satlantic ISUS v3)
Cast 001 - 020 (pre-calibration)
y = 36.487x - 1.361
R² = 0.9458

y = -18.685x2 + 57.757x - 2.2668

R² = 0.9623

Nitrate - ISUS 4sec ave voltage vs bottle NO3; Cast 021 - 075 (post-calibration) y = 27.587x - 4.030
R² = 0.9708
 
Fluorometer - linear & polynomial regressions

y=7.2233x-0.2435

R² = 0.7822

y = 9.4432x2 + 5.1519x - 0.1638

R² = 0.7863

  
 
  Secondary CTD O2 Sensor vs Bottle O2
MBARI-ISUS Voltage vs Bottle NO3 MBARI-ISUS Voltage vs Bottle NO3
Wetlabs ECO/FL Fluorometer Voltage vs Chlorophyll-a  
 
General notes: These are cast & preliminary CTD Processing Notes from 1601RL cruise

CalCOFI 1701RL General Cruise & Cast Notes:

Stations and Station Order: this cruise occupied 83 stations: 74 standard & 9 SCCOOS.  The station order was standard south-north, west-east-west but there was a 5-day severe weather delay after 75 stations. After the standard southern 75 station pattern was completed at 76.7 49.0, RV Reuben Lasker came back to San Diego (TAMT) for protection from a severe winter storm off most of California. Since the SIO personnel were to disembark in Monterey CA, they did so in San Diego.
The rosette was stripped of all bottles and CTD profile-only station casts were planned for the remaining stations (73.3 50.0, 73.3 55.0, 70.0 51.0, 70.0 55.0, 66.7 55.0, 66.7 50.0, 63.3 52.0, 63.3 55.0). CUFES egg survey transects were performed along a southward transect across station 60s after finishing northernmost sta 63.3 55.0. Refer to map for station order and track.

CTD General Notes and Problems:

The CTD configuration throughout the cruise was standard: Seabird 911+ with dual T, C, O2, & pumps;  Wetlabs C-Star 25cm transmissometer; Biospherical QSP200L PAR;  Datasonics/Benthos Altimeter; WET Labs ECO-AFL/FL; Seabird SBE 18 pH; Satlantic ISUS v3 (rebuilt & upgraded to firmware v3 prior to CalCOFI 1611SR; recalibrated by JRW/DGS this cruise after cast 020) & Wetlabs (custom) batteries. Please refer to the xmlcon files or cruise prospectus for additional info & metadata. There were no sensor failures this cruise but the ISUS-V6 user-polynomial estimated nitrate changed significantly after cast 020. The merged CTD+bottle data were processed accordingly - in two bins: casts 001-020 and 021-083. Please note that there are no bottle data for stations greater than 75 so use cruise-average corrected CTD data on casts 076-083.

The first leg of the cruise was relatively calm although there were 2 days of moderatley-rough weather and, later,the swell (ship-roll) on stations 071-075 picked up significantly before we sailed home. The 3500m deep casts on sta 90.90 or 80.90 were not performed - RV Lasker does not have enough wire. The transmissometer did not have the dropout problems of the previous cruise usually seen after the 3500m deep cast. The ISUS was mounted vertically, with optics pointing down worked well when powered but a few profiles were lost due to connector erosion (cast 060-062) or no power. pH sensor response was checked using standards before we sailed and pH response continues to be good. Coefficients were not changed from the last cruise and comparison to check DIC-pH samples are pending.

Logistics: CalCOFI 1701RL was CalCOFI's second cruise on NOAA R/V Reuben Lasker. We performed a similar Decemeber pre-holiday early load & setup since the departure was scheduled the day after the post-holiday return Jan 4th. Three days of loading and setup were performed at 10th Avenue Marine Terminal Dec 19 - 21. Final setup and securing were performed Jan 4 and 5. RV Lasker sailed with 12 scientists and 3 volunteers at 1300PST Jan 5th 2017. No acoustic calibration for the EK60 was performed.
RV Lasker's two-conductor termination was used with the Lasker's v2 deck unit and SIO-CalCOFI's primary PC (Windows 7 ASUS blade) for casts 001-016. Lasker's monitor, keyboard, and mouse were also interfaced with the SIO PC. RV Lasker's Seabird Remote Depth Readout Box was used, interfaced via serial port, not the deck unit. Lasker's NMEA GPS feed was also interfaced with SIO-CalCOFI's CTD PC using an additional serial port. Other than the primary data acquisition interface connected to com1, all other serial connections were USB-to-Serial adapters (carousel modem, NMEA GPS, remote depth box). SIO-CalCOFI's new (since 1611SR) LARS-designed 24-bottle rosette was used throughout the cruise.
During cast 017 (sta 93.120), the termination failed during the downcast at ~240m. ET Kim took ~3hr to reterminate the conductive wire after several trouble-shooting steps were performed to pinpoint that was the problem. Switching out of the Lasker's deck unit was one thing tried and SIO-CalCOFI's v2 deck unit was left connected to the SIO CTD computer after the successful deck test of the reterminated sea cable.
SIO-CalCOFI ran their own data network independent of the ship's network although some SIO computers did connect to the internet occasionally via RLWLAN wifi. Other interent devices, mainly smartphones and tablets, were connected all the time - one internet device connection per scientist at a time. This restriction helps with the limited bandwidth but eastward transects still have poor internet connectivity (prevalent on ALL ships we sail on except the recent Sally Ride CalCOFI).
The CESL sample log tablet and CELog event tablet performed reliably when hard-wired to the data network, as well as the Chl Van Flog fluorometer computer.
Portasal "Sally" had waterbath temperature controller issues and was replaced by the Lasker's backup Portasal (new, never used). Unfortunately, the Lasker's backup Portasal failed after one run so SIO Portasal "Harry" was transferred aboard at Dana Point. "Harry" worked fine for the remainder of the cruise and "Sally" was sent to Guildline for service & repair.
Note: SIO-CalCOFI CTD PC & deck unit were interfaced with R/V Reuben Lasker CTD monitor, keyboard, mouse, sea cable, & GPS for casts 017-083. KVM connections from the Lasker's rack-mounted CPU & deck unit were moved to our systems. Our remote-depth readout was not needed since RV Lasker's was installed and working, connected via com port. GPS was connected to a serial to USB FTDI dongle on the lower front usb port.

RV Reuben Lasker does not have a Knudsen 3 or 12kHz echosounder to find the bottom depth. The EK60 system worked well although a +5m offset for center board depth should be applied..

Cast Notes:

Cast 001 - no bottom valve was installed on bottle #5 so CTD was redeployed to get the 10m sample with bottle #8. Valve installed post-cast. Using RV Lasker's SBE 11 v2 deck unit, serial 11P62840-0892

Cast 002 - 003:  no console ops notes

Cast 004 - calm night

Cast 005 - computer time off by ~3mins vs GPS's UTC-8

Cast 006 - no issues or notes

Cast 007 - prodo station 93.40; NCOG samples collected

Cast 008 - no issues or notes; underway pH sample #12793

Cast 009 - bottle #20 may have been tripped in the air so #21 tripped at surface (start-of-cruise issue: deck ops were premature in retrieving the rosetter before the surface bottle was closed)

Cast 010 - DMW adjusted the computer clock to match GPS-8 time, minus 4mins (22:20 to 22:16)

Cast 011 - calm night; lots of pyrosomes & salps (sta 93.60)

Cast 012 - prodo station, a little early; nets first

Cast 013 - no issues

Cast 014 - wrong cast type initially selected, type II chl max at ~70m; underway pH #12794

Cast 015 - cold, alm, foggy night

Cast 016 - really early prodo, nets first but CTD cast started at 0723PST

Cast 017 - termination failure at ~240m on downcast, CTD brought back to surface. Deck unit is 0000000 - tried new fuses, SIO's deck unit - still 000000 so ET Kim called to reterminate. Wire disconnected from CTD and rosette - requested longer cable before the cable grip for a safety loop. (good thing early prodo was done last station). Original 1222PST cast time, now 1657PST - took ~4.5hrs for diagnosis and repair. Bottle #18 did not close so sent back down for 25m bottle (#21). Using SIO Deck Unit SBE11 v2 11plus 7-2009 (serial# 11P4306-0811).

Cast 018 - working fine; underway pH sample #12795

Cast 019 - cold, calm, rainy night - wind ~6.7kts

Cast 020 - prodo station; post-cast - ISUScom software used to download the ISUS data files and recalibrate the ISUS with DI (0um/l NO3).

Cast 021 - new ISUS calibration shifted baseline back to zero, low nitrates measurement improved significantly. Survey tech JGM at the console ops. Triggers #13 & #18 replaced

Cast 022 - no issues - underway pH #12808 taken

Cast 023 - calm night; downcast and upcast a bit different between 125-0m

Cast 024 - prodo station

Cast 025 - winch brake slipping during cast at 440m, 380m, and 170m - causes bottle to slowly get deeper before bottle closure so readjustment may be needed between writing numbers on console ops and tripping the bottle. Loopediting here would be a problem.

Cast 026 - no issues detected

Cast 027 - no winch slippage; calm night; bottle #21 mistripped - fired ok but no closure - serviced trigger post-cast; NCOGs filtered late - busy

Cast 028 - calm night, full moon; forgot to plug-in the ISUS battery so no NO3 profile. Bottom depth read by EK60 was 356m but CTD altimeter read ~45m to the bottom so after stopping at 345m, the CTD was adjusted to 380m using the altimeter. Bottle #21 was not needed but tripped to test cleaning effectiveness. ISUS battery replaced with backup, connector will be replaced later today.

Cast 029 - prodo station; nets first, new ISUS battery

Casts 030 - 033 - shallow coastal stations, no issues noted on console ops; underway pH sample #12842

Cast 034 - no issues noted on console ops but "wrong sta name in .hdr, rushed turn around" noted on other form, hexedit correction by JRW; underway pH #12842

Cast 035 - deep SCCOOS - ship hesitant to move closer to shore at night, bottom shoals up sharply here so 40+m bottom depth. Calm night, lots of debris in the water - loose eel grass, kelp. Old ISUS NO3 user polynomial deriving negative surface NO3 value - need new coefficients.

Cast 036 - calm night, light mist during the CTD cast and rain during the last half of sampling

Cast 037 - some rolls, seas picking up a little, cold but no rain; Santa Monica Basin cast to 725m. With rolls, taking a bit more time to trip at target depth

Cast 038 - prodo station

Cast 039 - San Nicolas Island station; one hour "delay' before commencing station ops

Cast 040 - no issues reported

Cast 041 - bottle #21 mistrip - O2 & chl samples taken from underway; trigger #21 serviced post-cast

cast 042 - moderate seas, cold night; tripped bottle #22 in case #21 didn't close again but #21 closed fine; bottle #21 JRW O2 test case

Cast 043 - prodo station

Cast 044 - bottle #22 closed as insurance for surface sample

Cast 045 - bottle #21 at surface closed fine; underway pH #12854

Cast 046 - Broad shallow (relatively speaking) chl max (40-90m) so Type I cast at sta 87.110! Rough seas, cold night - prepped the CTD upon arrival to station since seas were on the starboard (CTD) side. Down at 30 then 60m/min okay. CTD yo-yo'd between 40-30m on upcast since Jose didn't stop at 40m. bottle #22 closed as insurance but not needed

Cast 047 - prodo station

Cast 048 - no issues reported on console ops

Cast 049 - no issues reported on console ops; underway pH #12855 taken after station

Cast 050 - moderate seas, some BIG rolls, cold night; no real chl peak - chl from 0-50m mixed; bottle #1 & #21 did not close. Triggers replaced post-cast

Cast 051 - prodo station; triggers #1 & #21 replaced, new

Cast 052 - no issues reported on con ops

Cast 053 - extra bottle #22 closed but not needed

Cast 054 - calm night, shallow station; noted that bottom detecting transducer is on the centerboard so +5m should be added

Cast 055 - only ~109m to bottom here - went to 100m and winch crept to ~105m, had to be brought back up quickly to 100m before 1st trip (Had to wait ~10mins for Jose & Tom - original ETA was 0645 and we arrived early). Calm pre-sunrise shallow cast.

Cast 056 - forgot nav file; created one from 1st bottle mrk

Cast 057 - very shallow prodo station, only 4 prodo bottles; no issues

Cast 058 - sludgy oil slick floating by particularly during Bongo tow, waited several minutes for it to pass (GoPro videoed)

Cast 059 - no issues

Cast 060 - ISUS not plugged in but not noted on con ops, battery not replaced

Cast 061 - no ISUS data, battery not connecting; underway pH sample #12869

Cast 062 - some ISUS data, but only on upcast ~490-0m good profile

Cast 063 - calm cold night, some slow BIG rolls; new ISUS battery installed & working; bottle #15 accidental bottle-trip (clipboard touch spacebar!), no samples taken from extra 60m bottle. LTER had to renumber their bottles <60m.

Cast 064 - prodo station

Cast 065 - JLW noted "sensors acting strange - oxy & salt @20-40m @ start, okay on upcast (post-cast: replaying the cast, the pumps went off for some reason after first coming on - possible biofouling but odd it affected both sides; secondary downcast surface to 40m data are also bad)

Cast 066 - 2.5gals of pyrosomes in manta net (sta 80.90); underway pH #12915

Cast 067 - calm cold night; small pyrosomes at surface during CTD launch; chl max still shallow; some BIG rolls;

Cast 068 - prodo station; had to go back down to 110 from 100; underway pH #12926

Cast 069-070 - nothing reported

Cast 071 - Rough seas, BIG rolls, loopedit may be needed this cast; winch creeping aging, having to readjust target depth after stopping; O2 pretty "squirrely" but both agree; unusual profile features between 200-330m. Nice CTD landing.

Cast 072 - Rough seas, rough deployment - may have lost a bottle nozzle or two according to Jian (rosette was released at the rail w/o fending off the swing back inboard - hit the ship hard). Did not want to recover and repeat rough deployment so it was sent down as is. Loopedit recommended. Waited several minutes with CTD at surface after 2min 10m soak for ship to find a better heading. Smooth landing with 4 bodies (JRW helped inboard). One valve tip cracked but sample not compromised. Rosette frame cracked in two places (welds) and upper ring bent up slightly, causing stanchion to angle sideways slightly. It did not impede the closure of the bottle #17.

Cast 073 - ISUS dropped out during the downcast ~98-140m; fine on upcast

Cast 074-075 - no issues reported on last two stations although swell is picking up and we are heading south as soon as we finish the 075. Weather forecast is ominous (ride home proved it, it got extremely windy and rough during the transit to SD)

LEG II: 25-29 Jan 2017 (5-day weather delay)

Cast 076 - sta 73.50, no bottles - 30m profiles only, ISUS and all sensors still installed

Cast 077 - sta 73.55 515m profile only

Cast 078 - sta 70.51 285m profile only;

Cast 079 - sta 70.55 515m profile

Cast 080 - sta 67.55 - 515m cast but data archiving never started = no data; cast plots printed

Cast 081 - sta 67.50 200m profile

Cast 082 - sta 63.52 80m profile

Cast 083 - sta 63.55 290m profile

File notes:
Seasoft-generated asc-hdr files were not renamed to YY-YYLLLLSSSS_###d or u.asc & .hdr. Voltages were not relabled. This practice makes it difficult to reprocess & merge with bottle data if necessary.

Mislabels found and corrected:
 Cast 034 sta was mislabled in the .hdr, .hex. Hexedit-corrected post-cast

06JFeb2017

 

Go to top
Our website is protected by DMC Firewall!